OpenStudy (anonymous):

block 1 (mass=3.17kg) is connected by a massless string to block 2 (mass=2.37kg). An upward force with magnitude 107.5N is exerted on block 1 causing the two block system to accelerate upward. What is the magnitude of the tension in the string?

OpenStudy (anonymous):

The only thing i have been able to do so far is make free body diagrams and sum up the forces for both blocks in the y direction. When i solved for acceleration and plugged it into my other equation i got the following: F-T-mass#1*g=(mass#1-mass#2)(T-mass#2*g/mass#2) But now i have tension on both sides and i dont know how to solve for it

OpenStudy (anonymous):

I'm still a bit fuzzy on how the blocks and string look without a diagram, but I can work on your equation with you ^_^ "F-T-mass#1*g=(mass#1-mass#2)(T-mass#2*g/mass#2)" \[F-T-m_1*g = (m_1-m_2) \frac {T-m_2*g}{m_2}\] did i write your equation correctly?

OpenStudy (anonymous):

@me1987

OpenStudy (anonymous):

could you draw a diagram (not a free body) of how this problem looks?